The radius of a circular oil slick on the surface of a pond is increasing at the rate of 10 meters/min. At what rate is the circle's area changing when the radius of the oil slick is 20 m. ?


1
Expert's answer
2021-10-08T08:02:40-0400

Given

"\frac{dr}{dt}=10 m/min"

We have to find "\frac{dA}{dt}" at r=20 metrs

Now "A=\u03c0r^2"

"\frac{dA}{dt}=2πr\frac{dr}{dt}"

At r=20m

"\frac{dA}{dt}=2π \times 20 \times 10=400π \\

\frac{dA}{dt}=1256"


Answer 1256 m/min